Specifications for a part for a DVD player state that the part should weigh between 25.5 and 26.5 ounces. The process that produces the parts has a mean of 26.0 ounces and a standard deviation of .22 ounce. The distribution of output is normal. Use Table-A.

a.What percentage of parts will not meet the weight specs? (Round your "z" value and final answer to 2 decimal places. Omit the "%" sign in your response.)

Percentage of parts 2.32 % (this is correct)

b.Within what values will 99.74 percent of sample means of this process fall, if samples of n = 12 are taken and the process is in control (random)? (Round your answers to 2 decimal places.)
Answer #1

For this probelm, first we have to calculate Z value

Z=(X-\mu )/\sigma

Z=(26.5-26)/0.22

Z = 2.27

At z =2.27 we get probability which meet specification = 0.9884

Then percentage of parts will not meet the weight specs =1-0.9884= 0.0116

But we have two tails upper tail and lower tail.

% of parts will not meet the weight specs = 2*0.0116 = 0.0232 = 2.32%%

b)

Given probability = 99.74%

Z = 2.79

UCL = \bar{x}+Z\sigma /\sqrt{n}

uCL = 26+2.79*0.22/\sqrt{12} = 28.13

LCL = \bar{x}-Z\sigma /\sqrt{n}

LCL = 26-2.79*0.22/\sqrt{12}

LCL= 23.87